### v1

# Synthetic gain for electron-beam spectroscopy

Abstract
Electron-beam microscopy and spectroscopy featuring atomic-scale spatial resolution have become essential tools used daily in almost all branches of nanoscale science and technology. As a natural supercontinuum source of light, free electrons couple with phonons, plasmons, electron-hole pairs, inter- and intra-band transitions, and inner-shell ionization. The multiple excitations, intertwined with the intricate nature of nanostructured samples, present significant challenges in isolating specific spectral characteristics amidst complex experimental backgrounds. Here we introduce the approach of synthetic complex frequency waves to mitigate these challenges in free-electron–light interaction. The complex frequency waves, created through causality-informed coherent superposition of real-frequency waves induced by free electrons, offer virtual gain to offset material losses. This amplifies and enhances spectral features, as confirmed by our electron energy loss and cathodoluminescence measurements on multi-layer membranes, suspended nanoparticles, and film-coupled nanostructures. Strikingly, we reveal that our approach can retrieve resonance excitation completed buried underneath the zero-loss peak, substantially enhance the quality of hyperspectral imaging, and resolve entangled multiple-photon–electron events in their quantum interaction. Our findings indicate the versatile utility of complex frequency waves in various electron-beam spectroscopy and their promising diagnostic capabilities in free-electron quantum optics.
Free-electron–light interactions De Abajo (2010); Polman et al. (2019); Garcia de Abajo and Di Giulio (2021); Coenen and Haegel (2017); Talebi (2018); Rivera and Kaminer (2020); Shiloh et al. (2022); Roques-Carmes et al. (2023) lay foundations for advanced electron-beam microscopy and spectroscopy De Abajo (2010); Polman et al. (2019); Garcia de Abajo and Di Giulio (2021); Coenen and Haegel (2017); Hassan et al. (2017); Varkentina et al. (2022); Nabben et al. (2023); Bucher et al. (2024); Gaida et al. (2024), tunable radiation sources spanning from X-ray, ultraviolet, Terahertz to microwave regime Massuda et al. (2018); Roques-Carmes et al. (2019); Liu et al. (2017); Korbly et al. (2005); Shentcis et al. (2020); Yang et al. (2023); Gong et al. (2023); Huang et al. (2023); Roitman et al. (2024), integrated dielectric accelerators Sapra et al. (2020); Shiloh et al. (2021); Chlouba et al. (2023), free-electron lasers Pellegrini et al. (2016), and quantum optics advances based on free-electron–photon entanglement Bendana et al. (2011); Kfir (2019); Di Giulio et al. (2019); Di Giulio and de Abajo (2020); Ben Hayun et al. (2021); Di Giulio et al. (2021); Dahan et al. (2021); Kfir et al. (2021); Karnieli et al. (2021); Konečná et al. (2022); Baranes et al. (2022); Feist et al. (2022); Adiv et al. (2023); Synanidis et al. (2024); Rasmussen et al. (2024). Among these applications, electron energy-loss spectroscopy (EELS) and cathodoluminescence (CL) have arguably generated the most profound influence: they have evolved into daily-use methodology in physics, chemistry, and biomedicine because of their superior atomic-scale spatial resolution De Abajo (2010); Polman et al. (2019); Garcia de Abajo and Di Giulio (2021). Moreover, photon-induced near-field electron microscopy (PINEM) Barwick et al. (2009); Park et al. (2010); García de Abajo et al. (2010) combines ultrafast optics to electron microscopy and enables the measurement capability for free-electron–based ultrafast interaction dynamics on a sub-femtosecond time scale Feist et al. (2015); Priebe et al. (2017); Di Giulio et al. (2019); Vanacore et al. (2019); Synanidis et al. (2024). In a similar vein, electron energy gain spectroscopy (EEGS) aims to enhance spectral resolution by employing pulsed illumination with tunable frequency, reaching deep into the sub-microelectronvolt energy range Auad et al. (2023); Asenjo-Garcia and De Abajo (2013); de Abajo and Kociak (2008).
Despite these rapid advances, the naturally weak interaction between free electrons and photons, as evident by the small magnitude of the fine structure constant, leads to measurement challenges regularly encountered in electron microscopy and spectroscopy, where spectral features could be hard to resolve or even lost due to their limitations listed below. First, one has to deal with a low signal-to-noise ratio (SNR) Hofer et al. (2016); De Abajo (2010), especially when a high energy resolution is used in EELS. Meanwhile, CL probability suffers from low emission efficiency, resulting in a stringent SNR condition in CL spectroscopy. Second, because free-electron excitation naturally contains all frequency components, adjacent spectral features could be partially masked by or buried in the tails of their neighbors. This is particularly true for the zero-loss peak notorious for overshadowing low-energy excitations (covering the terahertz, infrared, and optical regimes), a unique limitation of electron spectroscopy that one needs to battle with compared with optical spectroscopy. Third, spontaneous events of multiple entangled photons de Abajo (2013); Feist et al. (2022); Adiv et al. (2023); Arend et al. (2024) are being heavily sought in free-electron quantum optics. Specifically, these multi-photon events obey the Poisson distribution, and thus, their probability becomes increasingly weak and calls for methods to confirm their experimental signatures.
One potential way to mitigate these limitations is to compensate for the intrinsic material losses such that spectral features become more pronounced. To this end, one may assemble real gain into samples to offset the loss Grgić et al. (2012); Fang et al. (2009); Ramakrishna and Pendry (2003); however, the addition of gain may induce instability and noise, and certain types of samples, such as those sensitive to heat, can be intrinsically incompatible with real gain. An alternative approach is introducing complex frequency wave (CFW) with virtual gain Archambault et al. (2012); Tetikol and Aksun (2020); Li et al. (2020); Kim et al. (2022). CFW with temporal attenuation has been shown to have the capability for compensating the loss in optical Tetikol and Aksun (2020); Li et al. (2020); Kim et al. (2022, 2023) and acoustic systems Gu et al. (2022); Kim et al. (2023). Recent advances have demonstrated a synthetic CFW approach featuring causality-informed integration of multiple real-frequency excitations. This synthetic approach has been demonstrated in enhancing superlens imaging Guan et al. (2023), phonon polariton propagation Guan et al. (2024), and molecular sensing Zeng et al. (2024). However, this progress all focuses on optical microscopy and spectroscopy, whereas their electron-beam counterpart remains unexplored.
This work introduces the CFW approach to free-electron–light interaction for enhancing electron-beam microscopy and spectroscopy with synthetic gain. After theoretically showcasing its effectiveness with canonical geometries, we demonstrate the general utility of this approach using three sets of measurements, including EELS measurements of a dielectric-metal-dielectric (DMD) multi-layer membrane, EELS measurement of single silver nanoparticles, and CL spectroscopy of film-coupled plasmonic nanospheres. We further apply our approach to electron energy gain spectroscopy and quantum-entangled free-electron multi-photon events. These efforts jointly show that the CFW approach enables the retrieval of weak hidden spectral features and the enhancement of the existing features, thereby substantially improving the capability for mode characterization, hyperspectral imaging, and quantum event detection in electron-beam microscopy and spectroscopy.
In EELS and CL, electron beams generate evanescent incident fields spanning all frequencies, interact with and get scattered by specimens, and lead to electron energy loss and radiation known as cathodoluminescence, as shown in Fig. 1a. Broadly speaking, it has been established that a passive scatterer in a vacuum is causal, meaning that it is analytic in the upper-half plane and satisfies the Kramers–Kronig (KK) relations Srivastava (2021). The free-electron manifestation of this principle can be mostly conveniently seen in small point particles, whose energy loss is simply proportional to the imaginary part of their polarizability Polman et al. (2019). For general situations, one could leverage a -matrix–based oscillator representation framework Zhang et al. (2023); Xie et al. (2024) (where the operator relates the polarization fields with incident fields Carminati and Schotland (2021); Chao et al. (2022)) and show that the energy loss is proportional to the imaginary part of the matrix (see Sec. S5), which is causal and passive simultaneously.
Based on this foundation, to augment the features in electron microscopy and spectroscopy, one potential mathematical method is to replace real frequency with a suitable complex frequency for loss compensation. This is equivalent to offset the imaginary part of material permittivity, e.g., transforming the permittivity of a Lorenz model into a purely real value, by the substitution of with . However, in this treatment, energy would diverge as time tends to negative infinity, indicating that the theoretically ideal CFW is unphysical. Practically, we can express a time-truncated CFW with the expression of , where is the original signal in complex form (e.g. electric fields), is the time-truncated signal, and is the Heaviside step function with at and otherwise. By leveraging the causality property of the -matrix, we can recover a loss probability with complex values based on the Kramers-Kronig relations, , where denotes the Cauchy principal value, and represents the real-frequency probability measured from experiments. Using Fourier transformation, a linear combination of real-frequency probability from experimental acquisition can be used to synthesize the complex-valued loss probability at complex frequencies :
and the real-valued probability associated with the complex frequency is given by .
Next, we theoretically demonstrate the synthetic CFW approach using two analytical scenarios as shown in Fig. 1b-g. The first example is an electron passing above a metal half-space at a distance of in vacuum, as shown in Fig. 1b, which allows for analytical treatment Lucas and Šunjić (1971); Echenique and Pendry (1975); Otto (1967); Garcia-Molina et al. (1985); Forstmann et al. (1991) (see Sec. S6A). When the separation is large at , the spectral signature at associated with surface plasmon launching, gets buried in the tail of the zero-loss peak, as seen in the blue curve before CFW processing in Fig. 1c. Aiming at restoring the SP peak, we apply the CFW processing based on Eq. (1). The post-CFW spectrum is shown by the red curve in Fig. 1c, where the previously buried surface plasmon mode can be seen clearly. It is noted that the red curve in Fig. 1c is shifted up by 0.5 relative to zero probability for better visualization (same treatment applied to all comparisons in other figures containing multiple spectra unless specified otherwise).
The second analytical example is an electron passing near a nanoparticle with an impact parameter , as shown in Fig. 1d. The CL spectra of this interaction permit analytical treatment De Abajo and Howie (1998); De Abajo (2010) which describes a summation of various Mie scattering channels labelled by integers (see Sec. S6B) . When the material loss of the nanoparticle is large, the raw spectrum (blue curve in Fig. 1e; for electrons at position B in Fig. 1d) displays a single-peak lineshape corresponding to the dipolar () mode. Nevertheless, both the dipolar and quadrupolar () modes can both be clearly seen after CFW processing (red in Fig. 1e). To further demonstrate the effectiveness of the CFW method, Fig. 1f displays a theoretical one-dimensional hyperspectral mode mapping imaging from Position A to C. The excitation of the dipolar mode predominates, whereas the quadrupolar mode can hardly be recognized. By applying CFW processing into the whole hyperspectral image, the previously hidden quadrupolar mode becomes vividly seen, leading to the occurrence of clear two-mode trajectories in the post-CFW mode mapping (Fig. 1g).
Motivated by such enticing performances as predicted in theory, we perform three measurements to showcase the capability of the CFW technique experimentally. These experiments are the energy loss measurement of a multilayer thin film and a suspended nanosphere, and the CL measurement of a film-coupled nanosphere.
We first conduct a bulk loss measurement of a dielectric-metal-dielectric (DMD) multilayer structure (Fig. 2a). The DMD structure consists of a thick Au layer sandwiched between two TiO2 thick layers, as shown in the cross-sectional TEM image (right of Fig. 2a). The originally measured bulk loss spectrum (blue curve in Fig. 2b) of 200-keV electrons displays three peaks at around , and , respectively. Extra spectral features emerge in the post-CFW EELS spectrum (red curve in Fig. 2b): the original three peaks become more pronounced, and more importantly, two new peaks at around and appear. Taken together, these five peaks are labeled as i to v according to their ascending frequencies. Furthermore, We conduct numerical simulations to calculate the energy loss probability (Fig. 2c) and induced field profiles (Fig. 2d) at those peaks. Peak i primarily arises from the transitions from the peak of the TiO2 valence band to the and levels of the titanium orbitals Launay et al. (2004), with a potential contribution from the bulk plasmon excitation in gold Motornyi et al. (2020). Besides, Peaks ii and iii are mostly associated with the electron-hole pair generation in TiO2 Vast et al. (2002); Mo and Ching (1995). Meanwhile, Peaks iv and v stem from the titanium semicore transitions Vast et al. (2002); Wang et al. (2009). The augmented spectral features after CFW processing are in good agreement with the numerical simulations (See Sec. S4) in Fig. 2c.
The next experiment focuses on suspended silver nanoparticles. Single silver nanoparticles of diameter around are suspended within holes of a transmission electron microscopy (TEM) grid (Fig. 3a). A 200-keV electron beam scans across a single nanoparticle, creating a scanning line with steps for mode mapping. The hyperspectral images along the scanning trajectory from A to C (see red line in Fig. 3a) are presented in Fig. 3b (top: before CFW; bottom: after CFW). The EELS intensity of the peaks varies along the scanned line in accordance with the coupling efficiency at which surface and bulk plasmons are launched under different impact parameters. The trajectories of the surface and bulk plasmon peaks become more pronounced after the CFW processing. To see this more clearly, Fig. 3c presents the raw EELS spectra measured at three representative positions (aloof positions A and C, and penetration position B) along the beam scanning line. At Position B, weak features at around and can be seen in the EELS spectrum without CFW, corresponding to the dipolar Mie surface plasmon resonance and bulk plasmons, respectively. Whereas at Positions A and C, spectral features cannot be recognized except for the zero-loss peak because the coupling between the electron and the nanoparticle gets weaker at larger separations. Strikingly, after the CFW processing (Fig. 3d), the dipolar Mie resonance peak emerges in the spectra of both Position A and C, where, in the meantime, the bulk plasmon peak remains absent because the interaction is in the aloof configuration. For Position B, the weak features of the surface and bulk plasmons are both substantially enhanced. In addition to EELS, in Sec. S7 we apply our approach to EEGS and exemplified its capability in resolving and characterizing closely neighbored resonances.
Next, we move to cathodoluminescence, another important type of spectroscopy that combines high spatial and spectral resolution. Fig. 4a illustrates the experimental structure—a film-coupled nanoantenna geometry consisting of single silver nanoparticles of diameter on a thick gold film, interacting with a electron beam. Again, we perform a hyperspectral CL imaging measurement along a scanning line across the nanoparticle (see red line in Fig. 4b). In the raw hyperspectral image as measured (Fig. 4c), the trajectory of only a single mode can be well recognized at , together with faint features of another mode around . Nevertheless, after CFW processing, four modes appear in total, and they are labeled as M1 to M4 with ascending frequency as shown in Fig. 4d. The spectra at representative positions A, B, and C before and after CFW processing are shown in Fig. 4e and f, respectively. To elucidate the origin of each mode, we perform numerical simulations, and the resulting spectra before and after CFW processing are shown in Fig. 4g, which corresponds well with the experimental data in Fig. 4e and f. We further elucidate the origin of the peaks using eigenmode calculations in Fig. 4h. Because of the sphere-film–coupled geometry, the gap plasmon modes and the Mie modes get hybridized. In the low-frequency region, M1 and M2 predominately display gap-plasmon features and are labeled by their quantized in-plane wavevectors and along the radial and azimuthal direction, respectively Yang et al. (2017). As the frequency increases, M3 and M4 primarily exhibit Mie characteristics and can be identified by the integer labels of spherical harmonics. Nevertheless, CL in this region becomes highly multi-channel in . Therefore, in Fig. 4, we display the eigen-fields of and for M3 and M4, respectively. These modes are the primary contributors to these two peaks according to the Mie CL calculation De Abajo (2010); De Abajo and Howie (1998).
Lastly, into the quantum interaction regime, we theoretically demonstrate that CFW can also be used for resolving experimental signatures of spontaneous free-electron multi-photon processes. This quantum interaction can be described by a model of Poisson statistics describing the spectral density of quantized photons that are equally separated by a single photon energy in the EELS spectrum Adiv et al. (2023); Feist et al. (2022) (see Sec. S8). Based on this model, we construct an EELS spectrum with a large quantum coupling strength such that multi-photon processes can spontaneously happen; the associated energy loss spectrum is shown by the blue curve of Fig. 5a. Despite the one-, two-, and three-photon peaks being evident, the four-photon peak (and even higher-order peaks) is poorly resolved even without noise. In practical EELS experiments, such a weak feature will be further contaminated by the addition of experimental noise. Nonetheless, after CFW processing, the four-photon peak becomes clearly recognizable. At the same time, all other lower-order photon peaks get enhanced (see red curves in Fig. 5a and the zoom-in image in Fig. 5b). This indicates that the CFW method could be particularly useful for restoring experimental signatures in free-electron quantum optics.
In conclusion, we have introduced an approach of complex frequency waves to provide virtual gain for electron microscopy and spectroscopy. In fact, our approach shares a similar rationale with electron energy gain spectroscopy de Abajo and Kociak (2008), yet differs in that EEGS relies on real gain from external laser pumping, whereas our CFW approach offers virtual gain through causality. We have showcased its capabilities in restoring spectroscopic features and improving hyperspectral imaging across various scenarios, including multi-layer membranes, suspended and film-coupled nanoparticles, and spontaneous multi-photon processes. Aside from the optical excitations mostly focused on in this work, our approach can be generally applied to a broad range of excitations in electron-beam spectroscopy, encompassing a spectrum ranging from low-energy phonons to high-energy ionization. In addition to the EELS and CL systems studied here, our CFW approach extends to optically pumped electron-beam systems like EEGS and PINEM. Going beyond electron-beam systems, this approach may be applicable to a large family of electron spectroscopy for condensed matters, such as X-ray photoelectron spectroscopy, angle-resolved photoemission spectroscopy, scanning tunneling spectroscopy, and so forth. The virtual gain provided by the complex frequency waves opens an avenue toward resolving subtle spectroscopic features in electron-beam microscopy and free-electron quantum optics.
Acknowledgments. We thank Germaine Arend, Frankie Y. F. Chan, Mathieu Kociak, Zemeng Lin, and Bengy T. T. Wong for experimental help and stimulating discussions.